# Turbulent dynamos in a collapsing cloud

### v1

@@ I Nonlinear dynamos
The rate at which magnetic field is amplified in the kinematic dynamo stage depends on both the free-fall time $t_{\rm ff}$ (thus on the initial cloud density) and the time when the collapse stops, i.e., on $a_{*}$ (if the dynamo is still kinematic at $t_{*}$), whereas the steady-state magnetic field strength depends only on $a_{*}$. We consider three cases where $t_{\text{ff}}$ is smaller, comparable to or larger than the time $t_{\text{nl}}$ when the standard dynamo becomes nonlinear and the magnetic field growth saturates to reach a quasi-steady state (note that $t_{\text{nl}}$ is different for the SSD and LSD): (i) $t_{\text{nl}}>t_{\rm ff}=50$, (ii) $t_{\text{nl}}\sim t_{\rm ff}=250~{}(\text{SSD}),\ 200~{}(\text{LSD})$, (iii) $t_{\text{nl}}<t_{\rm ff}=400$. In all cases, we stop the collapse at two different times, $t_{*}/t_{\text{ff}}=0.8$ and $0.95$ which correspond to different simulations with $a_{*}\simeq 0.5$ and $0.2$, respectively.
Figure 3: The rms supercomoving magnetic field strength compensated for the compression, $\tilde{B}_{\text{rms}}/B_{0}$, in the SSD, for various values of $t_{\rm ff}$. The solid and dotted curves represent two choices of the scale factor at which the collapse ends, $a_{*}=0.5$ and 0.2, respectively. The inset plot presents the physical $B_{\rm rms}$ focusing on the nonlinear stage. The green dots on the curves represent the point at which the collapse has been stopped. The blue curves correspond to $t_{\rm ff}<t_{\text{nl}}$ and the others to $t_{\rm ff}\gtrsim t_{\text{nl}}$.
Figure 4: As Fig. 3 but for the LSD.
Figures 3 and 4 show the evolution of magnetic field in the SSD and LSD, respectively. The main frames isolate the effect of the dynamo action on the supercomoving rms magnetic field strength (presenting $\tilde{B}_{\text{rms}}=a^{2}B_{\text{rms}}$) whereas the insets, focusing on the nonlinear regime, show how the physical rms magnetic field $B_{\text{rms}}$ grows due to both compression and dynamo action. As Eqs. (10) and (12) show, when $t_{\rm ff}$ is smaller (the cloud is denser), $\tilde{a}$ evolves faster and the magnetic field grows correspondingly faster in its kinematic stage. An important observation is the distinction in magnetic field evolution between the cases where $t_{\rm ff}<t_{\text{nl}}$ and $t_{\rm ff}\gtrsim t_{\text{nl}}$. To appreciate this, we note that in the stationary background the kinematic stage ends (and the magnetic field strength levels off) at $t_{\text{nl}}=250$ and 200 in the SSD and LSD, respectively, significantly later than in a collapsing cloud with $t_{\text{ff}}<t_{\text{nl}}$. As a result, denser clouds develop strong magnetic fields earlier. Furthermore, if the dynamos were inefficient in a stationary background, the collapse would help it to reach the end of the kinematic regime much faster. This kind of speed-up is less significant when $t_{\rm ff}\gtrsim t_{\text{nl}}$.
@@
The supercomoving magnetic field continues to grow in the nonlinear regime in line with Eq. (13) which shows that $\tilde{B}\propto\tilde{a}(\tilde{t})^{-1/2}$ arising from the weakening of the supercomoving Lorentz force. The physical field also grows due to the flux freezing as shown in Eq. (14). Despite a strong effect of the collapse on the kinematic dynamo, the saturated magnetic field strength is virtually independent of $t_{\rm ff}$ but is sensitive to the degree of compression, $a_{*}$. Figure 5 shows the saturated physical field strength as a function of $a_{*}$. In a remarkable agreement with Eq. (14), $B_{\rm rms}$ scales as $a_{*}^{-5/2}$. We note that the magnetic field $\mathbf{B}$ is random with vanishing mean value in the case of the SSD and has a significant mean component in the LSD. As shown in Fig. 5, both parts of the magnetic field in the LSD scale similarly with $a_{*}$.
Conclusions and discussion
@@
Dynamo action in a stationary environment leads to an exponentially fast amplification of a seed magnetic field. We have shown that in a collapsing cloud the dynamo leads to a faster, super-exponential growth of the magnetic field. This effect is stronger than just a superposition of compression and exponential growth: the variable $a^{2}B$, where the overall impact of compression is factored out, also grows super-exponentially because the instantaneous growth rate of the magnetic field increases in the course of the collapse. The analysis in this work has benefited from using a framework of supercomoving variables that allowed us to recover the MHD equations in the collapsing background in nearly their original form.
Apart from the effect on the rate of amplification, the collapse also enhances the steady-state strength of the magnetic field which scales with the scale factor $a$ as $B\propto a^{-5/2}$, as opposed to $B\propto a^{-2}$ for a frozen-in magnetic field. For a spherically-symmetric collapse, a frozen-in magnetic field increases with the gas density $\rho$ as $B\propto\rho^{2/3}$, whereas the dynamo action leads to $B\propto\rho^{5/6}$. Galaxies form from the intergalactic medium with a density enhancement by a factor of $10^{5}$$10^{6}$. Without any dynamo action, a frozen-in magnetic field would evolve from an initial value $B_{i}$ to $(\rho/\rho_{0})^{2/3}B_{i}\simeq 10^{4}B_{i}$. However, with a dynamo operating, the final field strength, in equipartition with the turbulent flow, is $10^{5}(4\pi\rho_{0}\tilde{v}_{0}^{2})^{1/2}$, where $(\rho/\rho_{0})^{5/6}\simeq 10^{5}$. This is significantly stronger than what is achieved by flux-freezing alone or by a standard dynamo without collapse. The contributions from the dynamo and flux-freezing (on top of the dynamo) can be separated, with the latter providing the dominant enhancement. We have discussed these effects using the standard turbulent dynamos which generate random (SSD) and large-scale (LSD) magnetic fields, but the analysis applies to any other fast dynamo.
The super-exponential growth of both random and mean magnetic fields in a collapsing turbulent cloud could significantly alter estimates of the age at which young galaxies develop observable magnetic fields. Current dynamo models which do not include this effect explain large-scale magnetic fields at the redshift $z\simeq 3$ [57, 58], whereas random magnetic fields can become significant even earlier. Ordered galactic magnetic fields have been reported at redshifts up to $z=2.6$ [59] and even $z=5.6$ [11]. The polarized dust emission detected can emerge from both anisotropic random magnetic fields (which naturally occur in spiral arms due to differential rotation [6]) and mean magnetic fields. Whatever is the nature of the galactic magnetic fields at high redshifts, its super-exponential amplification may be a crucial aspect of the theory. To illustrate the new opportunities, we note that, in the case $t_{\rm ff}\sim 0.1~{}t_{\text{nl}}$, where $t_{\text{nl}}$ marks field saturation at presumably an observable strength, the super-exponential growth over $15$ orders of magnitude reduces $t_{\text{nl}}$ of the standard dynamo by a factor of $10$ [51], significantly affecting our understanding of galactic magnetic fields at high redshifts [60]. Cosmological MHD simulations of galaxy formation suggest turbulent dynamo action [61, 62], but evidence for super-exponential growth—potentially a key dynamo signature in a collapsing environment—has yet to be presented.
Our results also apply to primordial star formation, leading to significantly stronger protostellar magnetic fields [41]. The average gas density in stars (of the order $10^{24}\,{\rm cm}^{-3}$) is much higher than that of molecular clouds (of order $10^{3}\,{\rm cm}^{-3}$) from which they form. Thus the compression factor is enormous. However, it is possible that the dynamo is inefficient in these environments due to a low Prandtl number or low ionization, leading to a very large $t_{\text{nl}}$. Even then the collapse can significantly enhance the dynamo, with the extent of the speed-up depending on how large the ratio $t_{\text{nl}}/t_{\rm ff}$ is. A more detailed calculation of both the speed-up timescales and saturated magnetic field strengths in early stars and galaxies will be presented in a future work.
@@
In general, local vortical motions in turbulent flows are amplified during any collapse (because of the specific angular momentum conservation), leading to faster eddy turnover time and larger dynamo growth rate. The results obtained here in the context of homologous collapse can be adapted to any mode of the collapse (e.g., the hierarchical collapse [48, 63]). In particular, the conclusion that dynamo action leads to a super-exponential amplification of the magnetic field applies to any collapsing system.
Acknowledgments
@@
We thank Godwin Martin and the members of the ICTS Plasma - Astrophysics group, particularly Vinay Kumar, for useful discussions. This research was supported by the Department of Atomic Energy, Government of India, under Project No. RTI4001. All simulations were performed on the Contra computing cluster at the International Centre for Theoretical Sciences, Tata Institute of Fundamental Research.
@@
II Dynamo equations and their solutions in supercomoving variables
@@
where $\text{D}/\text{D}t=\partial/\partial t+(\mathbf{u}\cdot\nabla),$ $\mathbf{u}$, $p$, and $\rho$ are the gas velocity, pressure and density, respectively, $\Phi$ is the gravitational potential, $\mathbf{B}$ is the magnetic field, and $\nu$ and $\eta$ are the kinematic viscosity and magnetic diffusivity, respectively. The equations are closed by relating the pressure and density with the help of an equation of state.
@@
II.1 The small-scale dynamo (SSD)
Kazantsev’s solution of the induction equation for the correlation function of the magnetic field in a random flow [49] remains one of the most often used models of the SSD in a stationary background [6]. The model assumes incompressible, statistically isotropic and homogeneous, mirror-symmetric, $\updelta$-correlated in time Gaussian random velocity field. Consider such an incompressible supercomoving velocity field $\tilde{\mathbf{v}}(\tilde{\mathbf{x}},\tilde{t})$ and its supercomoving correlation tensor $\tilde{T}_{ij}(\tilde{r})$,
$\tilde{v}_{i}(\tilde{\mathbf{x}},\tilde{t})\;\tilde{v}_{j}(\tilde{\mathbf{y}},\tilde{s})\right>=\tilde{T}_{ij}(\tilde{r})\,\updelta(\tilde{t}-\tilde{s})\,,$ (4)
where $\tilde{r}=|\tilde{\mathbf{x}}-\tilde{\mathbf{y}}|$ and we note that the velocity field is $\updelta$-correlated in both original and supercomoving variables. Since the continuity equation preserves its form when written in the supercomoving variables (Sect. 4.2 of Ref. [45]) and the gas density is assumed to be independent of position, ~$\tilde{\nabla}\cdot\tilde{\mathbf{v}}=0$ implies that the flow is incompressible in the physical variables as well, $\nabla\cdot\mathbf{v}=0$. The velocity correlation tensor can be represented in terms of the longitudinal and transverse parts, $\tilde{T}_{\text{L}}$ and $\tilde{T}_{\text{N}}$, respectively, as $\tilde{T}_{ij}=(\updelta_{ij}-\tilde{r}_{i}\tilde{r}_{j}/\tilde{r}^{2})\tilde{T}_{\text{N}}+\tilde{r}_{i}\tilde{r}_{j}\tilde{T}_{\text{L}}/\tilde{r}^{2}$, where $\tilde{T}_{\text{L}}$ and $\tilde{T}_{\text{N}}$ are related to each other because of the flow incompressibility, similarly to the relation (8) for the magnetic field correlators.
@@ II.1 The small-scale dynamo (SSD)
$a(t)v_{i}(t)\;a(s)v_{j}(s)\right>=\tilde{T}_{ij}(\tilde{r})\updelta\left(f(t)-f(s)\right).$ (5)
@@
$v_{i}(t)v_{j}(s)\right>=\tilde{T}_{ij}(\tilde{r})\updelta\left(t-s\right).$ (6)
@@
The magnetic field correlation tensor $\tilde{M}_{ij}(\tilde{r},\tilde{t})$ cannot be assumed to be $\updelta$-correlated in time [6], and it has the form
$\tilde{B}_{i}(\tilde{\mathbf{x}},\tilde{t})\;\tilde{B}_{j}(\tilde{\mathbf{y}},\tilde{t})\right>=\tilde{M}_{ij}(\tilde{r},\tilde{t}).$ (7)
Since $\nabla\cdot\tilde{\mathbf{B}}=0$, the magnetic field correlation tensor $\tilde{M}_{ij}$ can be written in terms of its longitudinal and transverse correlators $\tilde{M}_{\text{L}}$ and $\tilde{M}_{\text{N}}$, $\tilde{M}_{ij}=(\updelta_{ij}-\tilde{r}_{i}\tilde{r}_{j}/\tilde{r}^{2})\tilde{M}_{\text{N}}+\tilde{r}_{i}\tilde{r}_{j}\tilde{M}_{\text{L}}/\tilde{r}^{2}$, which are related as
@@
is responsible for the magnetic field amplification and $\tilde{\eta}_{\text{T}}=\eta+\tilde{T}_{\text{L}}(0)-\tilde{T}_{\text{L}}(\tilde{r})$ represents the dissipation due to the electric resistivity and turbulent diffusion. This equation can be reduced to a Schrödinger-type equation and using its solution in the WKBJ approximation leads to [64, 6]
$\tilde{\text{B}}_{\text{rms}}\propto\exp{\left(\tilde{\gamma}\tilde{t}\right)},\quad\tilde{\gamma}\simeq\tilde{v}_{0}/\tilde{l}_{0},$ (10)
where $\tilde{B}_{\text{rms}}$ is the rms magnetic field strength, and $\tilde{v}_{0}$ and $\tilde{l}_{0}$ are the integral speed and scale of the flow. Since $\tilde{v}_{0}$ and $\tilde{l}_{0}$ remain constant during the collapse, $\tilde{\gamma}=\text{const}.$
II.2 The large-scale dynamo (LSD)

### v2

@@ Abstract
The rate at which magnetic field is amplified in the kinematic dynamo stage depends on both the free-fall time $t_{\rm ff}$ (thus on the initial cloud density) and the time when the collapse stops, i.e., on $a_{*}$ (if the dynamo is still kinematic at $t_{*}$), whereas the steady-state magnetic field strength depends only on $a_{*}$. We consider three cases where $t_{\rm ff}$ differently compares with $t_{\text{nl}}$, the time at which standard dynamo becomes nonlinear: (i) $t_{\text{nl}}>t_{\rm ff}=50$, (ii) $t_{\text{nl}}\sim t_{\rm ff}=250~(\text{SSD}),\ 200~(\text{LSD})$, (iii) $t_{\text{nl}}<t_{\rm ff}=400$. In all cases, we stop the collapse at two different times, $t_{*}/t_{\text{ff}}=0.8$ and $0.95$ which correspond to different simulations with $a_{*}\simeq 0.5$ and $0.2$, respectively.
Figure 2: The rms supercomoving magnetic field compensated for the compression, $\tilde{B}_{\text{rms}}/B_{0}$, in the SSD, for various values of $t_{\rm ff}$ specified in the legend. The solid and dotted curves represent two choices of the scale factor at which the collapse ends, $a_{*}=0.5$ and 0.2, respectively, and the green circles mark the corresponding time. The inset presents the physical $B_{\rm rms}$ focusing on the nonlinear stage. The blue curves correspond to $t_{\rm ff}<t_{\text{nl}}$ and the others to $t_{\rm ff}\gtrsim t_{\text{nl}}$. The green curves represent the dynamo in a stationary background ($a=1$).
Figure 2 and Fig. 5 in the end matter show the evolution of magnetic field in the SSD and LSD, respectively. The main frames isolate the effect of the dynamo action on the supercomoving rms magnetic field strength (presenting $\tilde{B}_{\text{rms}}=a^{2}B_{\text{rms}}$) whereas the insets, focusing on the nonlinear regime, show how the physical rms magnetic field $B_{\text{rms}}$ grows due to both compression and dynamo action. As Eqs (8) and (9) show, when $t_{\rm ff}$ is smaller (the cloud is denser), $\tilde{a}$ declines rapidly, yielding faster magnetic field growth in the kinematic stage. An important observation is the distinction in magnetic field evolution between the cases where $t_{\rm ff}<t_{\text{nl}}$ and $t_{\rm ff}\gtrsim t_{\text{nl}}$. To appreciate this, we note that in the stationary background (green curve) the kinematic stage ends (and the magnetic field strength levels off) in the SSD and LSD, significantly later than in a collapsing cloud with $t_{\text{ff}}<t_{\text{nl}}$. As a result, denser clouds develop strong magnetic fields earlier. Furthermore, if the dynamos were inefficient in a stationary background, the collapse would help it to reach the end of the kinematic regime much faster. This kind of speed-up is less significant when $t_{\rm ff}\gtrsim t_{\text{nl}}$.
@@
The supercomoving magnetic field continues to grow in the nonlinear regime in line with Eq. (10) which shows that $\tilde{B}\propto\tilde{a}(\tilde{t})^{-1/2}$ arising from the weakening of supercomoving Lorentz force. The physical field also grows due to the flux freezing as shown in Eq. (11). Despite a strong effect of the collapse on the kinematic dynamo, the saturated magnetic field strength is virtually independent of $t_{\rm ff}$ but is sensitive to the degree of compression, $a_{*}$. Figure 3 shows the saturated physical field strength as a function of $a_{*}$. In a remarkable agreement with Eq. (11), $B_{\rm rms}$ scales as $a_{*}^{-5/2}$. We note that the magnetic field $\mathbf{B}$ is random with vanishing mean value in the case of the SSD and has a significant mean component in the LSD. As shown in Fig. 3, both parts of the magnetic field in the LSD scale similarly with $a_{*}$.
@@
Conclusions and discussion.— Dynamo action in a stationary environment exponentially amplify a seed magnetic field. We have shown that in a collapsing cloud the dynamo leads to a faster, super-exponential growth of the magnetic field. This effect is stronger than just a superposition of compression and exponential growth: the variable $a^{2}B$, where the overall impact of compression is factored out, also grows super-exponentially because the instantaneous growth rate of the magnetic field increases in the course of the collapse. The analysis in this work has benefited from using a framework of supercomoving variables that allowed us to recover the MHD equations in the collapsing background in nearly their original form.
@@
In general, local vortical motions in turbulent flows are amplified during the collapse (because of the specific angular momentum conservation), leading to a shorter eddy turnover time and larger dynamo growth rate as the collapse progresses. The results obtained here in the context of homologous collapse can be adapted to any mode of the collapse (e.g., the hierarchical collapse [49, 60]). In particular, the conclusion that dynamo action leads to a super-exponential amplification of the magnetic field applies to any collapsing system. As shown by us in [52], super-exponential magnetic field amplification is possible even when the turbulence is not continuously driven but rather decays. Motivated by our work and using the formalism developed here, [61] have addressed this possibility, and confirmed it via simulations.
Apart from enhancing the growth rate, collapse also increases the steady-state strength of the magnetic field. For a spherically-symmetric collapse, a frozen-in magnetic field increases with the gas density $\rho$ as $B\propto\rho^{2/3}$, whereas the dynamo action leads to $B\propto\rho^{5/6}$ in the continuously driven case. A steeper scaling with exponent between $2/3$ and $5/6$ is predicted even if the supercomoving turbulence is decaying [52]. Galaxies form from the intergalactic medium with a density enhancement by a factor of $10^{5}$$10^{6}$. Without any dynamo action, a frozen-in magnetic field would evolve from an initial value $B_{i}$ to $(\rho/\rho_{0})^{2/3}B_{i}\simeq 10^{4}B_{i}$. However, with a dynamo operating, the final field strength, in equipartition with the turbulent flow, is $10^{5}(4\pi\rho_{0}\tilde{v}_{0}^{2})^{1/2}$, where $(\rho/\rho_{0})^{5/6}\simeq 10^{5}$. This is significantly stronger than what is achieved by flux-freezing alone or by a standard dynamo without collapse. The contributions of the dynamo action and compression with flux-freezing (in addition to the dynamo amplification) can easily be separated in the framework of super-comoving coordinates. We discussed this for random (SSD) and large-scale (LSD) magnetic fields; however, the analysis applies to any fast dynamo.
The super-exponential growth of both random and mean magnetic fields in a collapsing turbulent cloud could significantly alter estimates of the age at which young galaxies develop observable magnetic fields. Current dynamo models which do not include this effect explain large-scale magnetic fields at the redshift $z\simeq 3$ [62, 63], whereas random magnetic fields can become significant even earlier. Ordered galactic magnetic fields have been reported at redshifts up to $z=2.6$ [64] and even $z=5.6$ [11]. The polarized dust emission detected can emerge from both anisotropic random magnetic fields (which naturally occur in spiral arms due to differential rotation [6]) and mean magnetic fields. Whatever is the nature of the galactic magnetic fields at high redshifts, its super-exponential amplification may be a crucial aspect of the theory. To illustrate the new opportunities, we note that, in the case $t_{\rm ff}\sim 0.1t_{\text{nl}}$, where $t_{\text{nl}}$ marks field saturation at presumably an observable strength, the super-exponential growth over $15$ orders of magnitude reduces $t_{\text{nl}}$ of the standard dynamo by a factor of $10$ [52], significantly affecting our understanding of galactic magnetic fields at high redshifts [65]. Cosmological MHD simulations of galaxy formation suggest turbulent dynamo action [66, 67], but evidence for super-exponential growth—potentially a key dynamo signature in a collapsing environment—has yet to be presented.
Our results also apply to primordial star formation. The efficient dynamo action resulting from the strong coupling between primordial gas and magnetic fields [68, 69, 70], combined with the enormous compression factor from the molecular cloud (density $n\sim 10^{3}\,{\rm cm}^{-3}$) to the protostellar core ($n\sim 10^{24}\,{\rm cm}^{-3}$), leads to significantly amplified magnetic fields in protostars [26]. However, it is also possible that the dynamo is inefficient in these environments due to a low Prandtl number or weak turbulence, leading to a very large $t_{\text{nl}}$. Even then the collapse can significantly enhance the dynamo, with the extent of the speed-up depending on how large the ratio $t_{\text{nl}}/t_{\rm ff}$ is. A more detailed calculation of both the speed-up timescales and saturated magnetic field strengths in early stars and galaxies will be presented in a future work.
Our analysis focuses on homogeneous, incompressible supercomoving turbulence in a collapsing cloud, both decaying and maintained by an explicit forcing, using a formalism which can be extended to compressible flows. The form of the forcing used reproduces the amplification of random flows in the gravity-driven turbulence [30] although the effects of the density inhomogeneities remain to be explored.
@@
Acknowledgments.— We thank Godwin Martin and the members of the ICTS plasma astrophysics group, particularly Vinay Kumar, for useful discussions. This research was supported by the Department of Atomic Energy, Government of India, under Project No. RTI4001. All simulations were performed on the Contra computing cluster at the International Centre for Theoretical Sciences, Tata Institute of Fundamental Research. We thank the referees for very positive and helpful comments that have led to several improvements in the paper, in particular, regarding the dynamo action in decaying turbulence. M.I.P. acknowledges the warm hospitality of Inter-University Centre for Astronomy and Astrophysics (IUCAA).
Data availability.— The RMS magnetic field data that support the findings of this article are openly available [71]. The spectral data is storage heavy and can be provided upon reasonable request.
@@
I Dynamo equations and their solutions in supercomoving variables
@@
where $\text{D}/\text{D}t=\partial/\partial t+(\mathbf{u}\cdot\nabla),$ $\mathbf{u}$, $p$, and $\rho$ are the gas velocity, pressure and density, respectively, $\Phi$ is the gravitational potential, $\mathbf{B}$ is the magnetic field, $\nu$ and $\eta$ are the kinematic viscosity and magnetic diffusivity, respectively. The equations are closed by relating the pressure and density with the help of an equation of state.
@@
I.1 The small-scale dynamo (SSD)
[…]
@@ I.1 The small-scale dynamo (SSD)
$\left<a(t)v_{i}(t)\;a(s)v_{j}(s)\right>=\tilde{T}_{ij}(\tilde{r})\updelta\left(f(t)-f(s)\right).$ (17)
@@
$\left<v_{i}(t)v_{j}(s)\right>=\tilde{T}_{ij}(\tilde{r})\updelta\left(t-s\right).$ (18)
@@
[…]
Since $\tilde{\nabla}\cdot\tilde{\mathbf{B}}=0$, the magnetic field correlation tensor $\tilde{M}_{ij}$ can be written in terms of its longitudinal and transverse correlators $\tilde{M}_{\text{L}}$ and $\tilde{M}_{\text{N}}$, $\tilde{M}_{ij}=(\updelta_{ij}-\tilde{r}_{i}\tilde{r}_{j}/\tilde{r}^{2})\tilde{M}_{\text{N}}+\tilde{r}_{i}\tilde{r}_{j}\tilde{M}_{\text{L}}/\tilde{r}^{2}$, which are related as
@@
[…]
$\tilde{\text{B}}_{\text{rms}}\propto\exp{\left(\tilde{\gamma}\tilde{t}\right)},\quad\tilde{\gamma}\simeq{\frac{\tilde{v}_{\nu}}{\tilde{\ell}_{\nu}}=}\frac{\tilde{v}_{0}}{\tilde{l}_{0}}{\tilde{\text{Re}}^{1/2}},$ (22)
where $\tilde{B}_{\text{rms}}$ is the rms magnetic field strength, $\ell_{\nu}$, $v_{\nu}$ are the viscous eddy scale, velocity, and ${{\tilde{\text{Re}}}}={\tilde{v}_{0}\tilde{l}_{0}/\nu}$ with $\tilde{v}_{0}$ and $\tilde{l}_{0}$ are the integral speed and scale of the flow. Since $\tilde{v}_{0}$ and $\tilde{l}_{0}$ remain constant during the collapse, $\tilde{\gamma}=\text{const}.$ The initial kinematic stage continues until the magnetic energy reaches equipartition with the kinetic energy of the viscous eddies. After that the growth rate is governed by the next larger eddy and it continues until magnetic energy reaches equipartition with the integral scale over a time scale of the largest eddy turnover time.
While the $\delta$-correlation assumption simplifies our analysis, it is not essential. In standard dynamos, finite correlation times modify growth rates to some degree but preserve exponential growth [53, 54]; we thus expect the key finding of super-exponential growth to remain robust.
I.2 The large-scale dynamo (LSD)